OpenStudy (anonymous):

******medals and fans****Janette went to Tire Mart and paid $122.57 per tire for 4 tires and a flat rate fee for installation for a total of $555.28 before taxes. Monique went to Fast Wheels and paid $136.71 per tire for 4 tires and a flat rate fee for installation for a total of $585.34 before taxes. The difference in the amount that they paid in installation charges is ?$.

OpenStudy (anonymous):

Symbolize and Solve Equations

OpenStudy (anonymous):

Need help can someone help me and not just sit in this room and look?

OpenStudy (anonymous):

Sorry it wasn't letting me respond. Tire Mart: $122.57(4) + x = $555.28 $490.28 + x = $555.28 x = $555.28 - $490.28 x = $65 instillation fee Fast Wheels: $136.71(4) + x = $585.34 $546.84 + x = $585.34 x = $585.84 - $549.84 x = $36 instillation fee Fast wheels' instillation fee is $29 cheaper than Tire Marts